OpenStudy (anonymous):

Simplify the following DIVISION of radicals expression. Make sure you rationalize the denominator if needed.

OpenStudy (anonymous):

\[\sqrt{9/2}\]

OpenStudy (anonymous):

\[3/\sqrt{2}\]

OpenStudy (anonymous):

Then rationalize denominator

OpenStudy (anonymous):

multiply top and bottom by \[\sqrt{2}\]

OpenStudy (anonymous):

you will get \[3\sqrt{2}/2\]